OpenStudy (anonymous):

solve the equation Hint (look before you leap) x^2-(sqrt2+sqrt5)x+ sqrt10=0

OpenStudy (mertsj):

(x-sqrt2)(x-sqrt5)=0 x=sqrt2, sqrt5

OpenStudy (mertsj):

\[(x-\sqrt{2})(x-\sqrt{5})=0\]

OpenStudy (saifoo.khan):

\[x^2 - x \sqrt2 - x \sqrt5 + \sqrt{10} = 0\] \[x(x - \sqrt2) - \sqrt5(x - \sqrt2) = 0\]

OpenStudy (saifoo.khan):

Sorry did a typo.

OpenStudy (anonymous):

I came up with x=sqrt2 and x=sqrt 5

OpenStudy (saifoo.khan):

It's correct. :)

OpenStudy (anonymous):

cool thanks